The driver of a car sets the cruise control and ties the steering wheel so that the car travels at a uniform speed of 15 m/s in a circle with a diameter of 130 m . Through what angular distance does the car move in 4.10 min ? Express your answer using two significant figures. What arc length does it travel in this time? Express your answer using two significant figures.
Answer #1

radius R = 65m

linear spedd v = 15 m/s

angular speed is "linear speed / radius"
here it gives 15/65 = 0.23 rad/s
1) angular distance = angular speed*time

                               = 0.23*4.1*60 = 57.77 rad


2) arc length = radius*angle swept

   = 65*(58.5*2pi) = 3755.1 m